THE INTENSITY THEORY OF OVERTONES AND COMPOUND TONES IN RAMAN SPECTRA

\underline{V.Ye. POGORELOV}, Department of Physics, Kiev Taras Shevchenko’s University, Glushkov pr. 6, Kiev 252127, Ukraine.

Using the intensity theory of fundamental tones (impulse notion for description of light interaction with matter) it have been made more precise expressions for intensities of overtones and compound tones. It have been established theoretically and experimentally that even at absent of exact resonance the intensities of overtones and compound tones depend on exiting light frequencies more abruptly than for corresponding fundamental tones.